OpenStudy (anonymous):

Hey guys. I need some help. I keep getting the answer to be 13 and 11 but it doesn't add up to 48. The perimeter of a rectangular writing pad is 48 cm. The width is 2 cm less than the length. Find the length and the width. Help?

OpenStudy (anonymous):

w=L-2 (given) P=2(L+W) did you solve this one?

OpenStudy (anonymous):

Yes I did

OpenStudy (amistre64):

define perimeter please

OpenStudy (anonymous):

P=2(L+W)

OpenStudy (amistre64):

Now what is 2(13+11) ?

OpenStudy (anonymous):

actually your answer if right.... when you check, you forgot the P=2(13+11)=48

OpenStudy (anonymous):

Thank you inik!

OpenStudy (anonymous):

you did good! :)